OpenStudy (anonymous):

solve the radical equation. sqrt{x}+sqrt{x+5}=1

OpenStudy (anonymous):

\[\sqrt{x}^2=(1-\sqrt{x+5})^2\]

OpenStudy (anonymous):

\[x=x-2\sqrt{x-5}+6\]

OpenStudy (anonymous):

\[3=\sqrt{x-5}\]

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

x=14